OpenStudy (anonymous):

Hey is 2 times 5 10?

OpenStudy (anonymous):

That will be one half of (180 minus the interior angle of a regular hexagon). A regular hexagon can be divided into 4 triangles, so the sum of the interior angles of a hexagon is: 4 * 180 = 720 The interior angle of a regular hexagon is 720/6 So, x = (1/2)(180 - 720/6)

OpenStudy (anonymous):

All good now, @victoria650 ?

OpenStudy (anonymous):

x = (1/2)(180 - 720/6) -> x = (1/2)(180 - 120) x = (1/2)(60) -> x = 30

OpenStudy (anonymous):

You are very welcome! So appreciative! @victoria650